how do you solve: (1-4i)/(2+3i)

i indicates a number that is imaginary ( sqrt of -1)

thanks so much

Answer by Alan3354(69443)

how do you solve: (1-4i)/(2+3i)
-------
You don't solve it. There's no variable to solve for.
The goal is to rationalize the denominator, ie, get the i out of it.
---------
(1-4i)/(2+3i)
Multiply the NUM and DEN by the conjugate of the DEN, 2-3i
= (1-4i)(2-3i)/((2+3i)(2-3i))
= (2 - 11i +12i^2)/(4 - 9i^2)
= (2 - 11i - 12)/(4 + 9)
= (-10 - 11i)/13
or -10/13 - i(11/13)